The average rate of emission radioactive particles from a source was measured over a long period, and found to be 10 particles per unit time. After an experimental treatment had been applied to the source, a further sample was examined and emitted 17 particles in unit time. Test at 5% the null hypothesis that the rate of emissions is unchanged.
1
Expert's answer
2020-04-27T18:55:39-0400

The number of emitted particles has Poisson distribution.

"P\{\xi=k\}=\frac{\lambda^k}{k!}e^{-\lambda}, k=0,1,\ldots"

We have "\lambda_1=10, \lambda_2=17."

We will take samples of sizes "n_1=20" and "n_2=20" from the first and the second populations.

"P\{\xi=20\}=\frac{10^{20}}{20!}e^{-10}\approx 0.0019.\\
P\{\xi=20\}=\frac{17^{20}}{20!}e^{-17}\approx 0.0692."

We get relative frequencies:

"\omega_1=0.0019\\
\omega_2=0.0692"

"H_0: p_1=p_2; H_1: p_1\neq p_2\\
\text{where } p_i \text{ is probability that event A (the particle emits)}\\
\text{will happen in the i-th population}."

We will use the following random variable:

"u=\frac{\omega_1-\omega_2}{\sqrt{(\omega_1+\omega_2)(1-(\omega_1+\omega_2))(\frac{1}{n_1}+\frac{1}{n_2}})}\\
u_{obs}\approx -0.8281\\
\Phi(u_{cr})=\frac{1-\alpha}{2}=0.475\\
u_{cr}=1.96\\
(-\infty,-1.96)\cup (1.96,\infty)\text{ --- critical region}.\\
u_{obs}\text{ does not fall into the critical region. So we accept } H_0."

The rate of emission did not change after an experimental treatment.